OpenStudy (anonymous):

Part 1: Graphing: Is when you want a visual display of the equations or when you want to estimate a solution. Substitution: Is when one equation is already solved for one of the variables or when it is easy to solve for one of the variables. Elimination: Is when the coefficients of one variable are the same or opposites or when it is not convenient to use graphing or substitution.

OpenStudy (anonymous):

Part 2: I used (-y + 3x = 6) (y = -6x + 12) So how I did this was I did (-6 + 12) + 3x -6. So -6 + 12 is 6. Then I did 6x + 3x -6 - 12. So 6x + 3x is 9x, and -6 -12 is -18. Then I did 9x divided by 9x and that canceled out. Then I did -18 divided by 9 is -2. So x is -2. Then I did -6(2) + 12 and -6 times 2 is -12. Then I did -12 + 12 which equals to 0. So y = 0. So the solution to this problem is (0,-2)